### v2

# LiDDA: Data Driven Attribution at LinkedIn

Abstract.
Data Driven Attribution, which assigns conversion credits to marketing interactions based on causal patterns learned from data, is the foundation of modern marketing intelligence and vital to any marketing businesses and advertising platform. In this paper, we introduce a unified transformer-based attribution approach that can handle member-level data, aggregate-level data, and integration of external macro factors. We detail the large scale implementation of the approach at LinkedIn, showcasing significant impact. We also share learnings and insights which are broadly applicable to the marketing and ad tech fields.
Multi-Touch Attribution, Media Mix Modeling, Attention Mechanism, Computational Marketing, Deep Learning
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: XXX; XXX, 2025; XXXisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Artificial intelligenceccs: Applied computing Marketingccs: Information systems Online advertisingccs: Mathematics of computing Probability and statistics
1. Introduction
Marketers rely on attribution measurement to understand the impact of marketing campaigns on customers’ journeys and decisions. Accurate attribution enables marketers to better optimize marketing spend and strategy to maximize their return on investment (ROI). In this paper, we propose a framework for large scale end-to-end attribution system and details the implementation at LinkedIn. LinkedIn Data Driven Attribution (LiDDA) leverages a combination of bottom-up and top-down modeling approaches in order to build a more accurate attribution platform to measure campaign performance and ROI. The system was originally developed for our internal Go-To-Market (GTM) Marketing and has since expanded to support external applications for LinkedIn customers.
1.1. LinkedIn Use Case
GTM Marketing is responsible for managing digital marketing campaigns for various LinkedIn products (e.g., Recruiter Lite, Sales Navigator), targeting both acquisition and retention objectives, where marketing audiences include both consumers and buyers. Consumer marketing focuses on the individual as a member such as member account signups. Buyer marketing focuses on the company-level conversions such as the use of LinkedIn Ads platform ads or Talent Solutions job posts.
Both segments are targeted using a combination of owned and external channels. Owned channels include Email and the LinkedIn Ads Platform (Feed Ads) where we own the platform used and have greater visibility to individual engagements. External channels refer to other online platforms such as Paid Search, Social Media, and Streaming. These channels do not provide individual level data, and we are often limited to aggregate information such as the total number of impressions in a time period.
One major challenge is lack of a unified view of the member journey across multiple channels. Existing media-mix-modeling (MMM) approaches provide channel-level visibility; however they lack detailed campaign breakouts and journey information. Last touch attribution (LTA) was the default standard; yet it biased results to bottom of funnel (BOFU) campaigns and omitted channels without individual identifiers.
Through our work on LiDDA, we are able to build an approach which combines signals using both top-down and bottom-up information. This enables our marketing team to utilize attribution across different levels of granularity while still maintaining consistency. In the next section we further discuss the differences in top-down and bottom-up attribution approaches.
1.2. Overview of Related Work
Data-driven attribution (DDA) aims to measure the contribution of online campaigns at the user level using a bottom-up approach. On the other hand, Media-Mix modeling (MMM) considers both online and offline channels along with seasonality using a top-down (aggregate time-series) approach.
1.2.1. Media-Mix Modeling
Media mix models (MMM) uses aggregated time-series data to model conversions as a function of online and offline channels as well as macro-economic factors and seasonality. MMM have been widely used since the 1960s (Borden, 1964; McCarthy, 1978), and provide a wholistic view across all channels and external factors. On the other hand, they come with different challenges (Chan and Perry, 2017). One of the shortcomings is that they do not have visibility to analyze data at the user level.
1.2.2. Rules Based Attribution
Traditional attribution models commonly use rule-based attribution (RBA) approaches such as first-touch, last-touch, or time decay(Gaur and Bharti, 2020). These approaches assign conversion credit based on simple, predetermined rules at user level. RBA methods are easy to understand and implement however, they are often based on limited assumptions and introduces biases.
1.2.3. Data-Driven Attribution
Data-driven attribution (DDA) addresses limitation of RBA by leveraging machine learning and statistical techniques to allocate conversion credit learned from Data. Different modeling approaches such as Logistic regression (Shao and Li, 2011), Bayesian (A. Li, 2014; Lizhen Xu, 2014) and Survival Theory (Zhang et al., 2014) have been used. Deep learning approaches such as CAMTA (Kumar et al., 2020), DNAMTA(li et al., 2018), and CausalMTA (Yao et al., 2022) have demonstrated improved performance with these classes of models.
1.3. Exploration on Existing Approaches
Ideally, credit is assigned according to its relative casual contribution. In practice, obtaining the ground-truth data for causal attribution is challenging. First, members are potentially exposed to dozens of overlapping treatments across different channels both on and off of our platform. Second, simulating realistic counterfactual journeys is further complicated by requiring touchpoint dimensions (eg. campaign metadata, timestamps) and the presence of sequential biases where subsequent engagements are influenced by prior interactions.
Given these challenges, we explored approaches to approximate the actual attribution for computational efficiency at scale. We first evaluated the Logistic Regression approach by Shao & Li (Shao and Li, 2011). Although straightforward, it requires making a separate prediction for every timestep in the sequence and using the change in outcome probabilities as credit. The model also lacks inherent ability to model sequences. We additionally observed a bias that early touches had larger incremental changes before plateauing as paths grew in length.
We also evaluated DNAMTA (li et al., 2018) which uses attention over LSTM however we found that the LSTM layer yielded end of path biases. This is likely because those hidden states had accumulated a greater amount of information useful for the classification task. We recognized that the use of transformer type of attention offered both single-pass inference and an explanability mechanism without requiring the LSTM component. Also, we realized that existing DDA models lacks consideration of missing data, whether from inherent availability gaps or privacy constraints, and do not adequately balance internal versus external channels or integrate with macro‑level MMM frameworks.
1.4. Our contributions
As far as we know, LiDDA is the first industry deployed large scale DDA system that leverage transformer-type attention mechanism and can handle member-level, aggregate-level and integration of external macro factors. In particular, the system is highlighted by: (i) Transformer Architecture: our use of transformer type of attention offered both single-pass inference and an explanability mechanis; (ii) Temporal-aware attention: we discretize time gaps between campaign touchpoints and learn positional embeddings—plus a special day‑of‑week embedding—to capture both sequence order and seasonal effects; (iii) Privacy‑driven data imputation: to correct for missing paid‑media signals (e.g. due to GDPR/CCPA restrictions), we probabilistically inject aggregated impressions back into user paths; (iv) Unified DDA&MMM outputs: we reconcile discrepancies between independent DDA and macro‑level MMM by training-time calibration; (v) End‑to‑end solution: we present our modeling innovations, implementation details, comprehensive evaluation and insights from launching to both GTM marketing and Ads platform in the paper.
2. LiDDA System
2.1. System Overview
LiDDA is our modeled attribution platform (Figure  1) which utilizes an attention-based modeling approach. Attention models are well suited for this task since they retain sequential information about the buyer journey and also enable us to incorporate additional features about the member, their company, and campaign information.
Overall, the model is structured as a binary classification task to predict whether a buyer journey will result in a conversion $P(Y|E_{M},E_{C},S),$ where $Y$ denotes the binary conversion outcome, $E_{M}$ denotes the member embedding with length $D_{M}$, $E_{C}$ denotes the (member’s) company embedding with length $D_{C}$, and $S$ denotes the sequence of marketing touchpoints on the member with length $N$. $E_{M}$ is a member embedding derived from LinkedIn data such as skills, titles, etc. that generally capture latent features of that member. Likewise, $E_{C}$ is a company embedding derived from LinkedIn data such as industry, roles, etc.
From the trained model, we output the aggregated attention weight matrix from the touchpoint sequence $S$. The attention weights are normalized such that they add up to 1.0, i.e., $\hat{\alpha}_{i}=\frac{\alpha_{i}}{\sum_{j=1}^{n}\alpha_{j}},$ and the normalized weights are interpreted as percent contribution of that touchpoint to the final conversion event. We further discuss attention in Section 2.4.
The attention weights represent the relative credits within the path, however they still need to be rescaled for overall impact. We use a separate media-mix model (MMM) to adjust the MTA credits at the channel level to account for overall incrementality and channels where individual exposure data does not exist or is not available. This is discussed further in Section 2.7.
Figure 1. High level diagram showing the end-to-end components of our attribution modeling framework
[…]
2.2. Path Creation
2.2.1. Sequence Representation
The touchpoint sequence $S$ is a time series of touchpoints $t\in\mathcal{T}$. $K=|\mathcal{T}|$ is the distinct number of possible (channel, action) pairs with 0 reserved for padding. An example of a (channel, action) could be $(EMAIL,OPEN)$ or $(AD,CLICK)$. The sequence $S$ is limited to the last $N$ touches such that paths with a length greater than $N$ have their earliest engagements truncated. The choice of touchpoints included are context-specific and could represent marketing activities or other attributable steps in a defined journey.
During modeling, each touchpoint is represented as an integer and represented as a learned embedding $E_{T}$ where $E_{T}\in\mathbb{R}^{T\times{\text{dim}}}$ which captures the latent information about the touchpoint. The embeddings are concatenated along the time dimension such that each path $S$ is represented as a $N\times d$ tensor.
2.2.2. Touchpoint Representation
An additional feature, $E_{MCID}$, is concatenated to the learned $E_{T}$ at each timestep. The embedding $E_{MCID}\in\mathbb{R}^{MCID\times{\text{dim}}}$ is a representation of our marketing campaign which captures different aspects of the campaign such as its metadata or content. We obtain these embeddings from an upstream model which extracts a text description of the campaign and generates an embedded representation using an internal LLM fine-tuned on a LinkedIn corpus. This approach allows flexibility in combining structured and unstructured data consistently across domains.
2.2.3. Offline Processing
Our data pipeline process member events into marketing paths and augment the necessary features for modeling. We store these paths where they serve a dual purpose as model inputs and for use by marketing teams in post-modeling reporting. For model inference, we only define paths with respect to conversion events. The conversion event is the unique path identifier and anchor for all preceding touchpoints. During model training, we also generate a sample of non-converting paths with respect to the current date to provide negative labels to our model. We utilize immutable conversion and touchpoint identifiers that allow downstream partner teams to easily connect attribution results to other data for further analysis and reporting.
2.3. Positional Representations
2.3.1. Positional Encodings
Attention models do not inherently model the order of touches in the paths so additional encodings are needed to capture the relative positions of each touch. Sinusoidal positional encodings are commonly used for temporal ordering as in the original attention design (Vaswani et al., 2017). These are implemented as: $PE_{(pos,2i)}=sin(pos/10000^{2i/d_{model}})$, $PE_{(pos,2i+1)}=cos(pos/10000^{2i/d_{model}})$
Because our touchpoint embedding $E_{T}$ is at the granularity of (channel, action), the number of possible combination is far lower than typical NLP vocabulary sizes. Therefore we use a substantially smaller embedding dimension for our data. We employ Time Absolute Position Encodings (tAPE) where the positional encoding above is modified to adjust for the lower dimensional representation(Foumani et al., 2023). The tAPE encodings are defined as: $\omega_{k}=10000^{-2k/d_{model}}$, $\omega_{k}^{new}=\frac{\omega_{k}\times d_{model}}{L}$.
2.3.2. Positional Embeddings
Positional encodings are not sufficient for our domain where marketing engagements are generated with irregular patterns and not across fixed intervals. For example, the time between an impression and the following click may be a few seconds while the time between two impressions could be days. We capture these irregular differences by representing the time difference between interactions as a discrete interval. For LinkedIn use cases, we use days as the interval unit since the look-back-window is limited to a few months based on business requirements.
Each touchpoint has a time interval relative to the final conversion state in the path, encoded by days as $D$ with a corresponding learnt embedding $E_{D}$. Due to the strong seasonality of our marketing data, we also learn embeddings $E_{DOW}$ corresponding to DOW encoding of each touchpoints.
We combine all the embeddings together for each touchpoint as: $concat(E_{MCID},E_{T})+E_{D}+{E_{DOW}}+tAPE$.
2.4. Attention Mechanism
We explored different methodologies for performing data driven attribution in our model as discussed in Section 1.2.3 and Section 1.3. In addition to their efficiency, models incorporating attention layers such as DNAMTA and CAMTA (li et al., 2018; Kumar et al., 2020), outperformed other model designs in this domain. Attention was originally proposed for NLP tasks (Vaswani et al., 2017), but later adopted in many other domains. Formally, attention is defined as:
(1) $\mathrm{Attention}(Q,K,V)=\mathrm{softmax}(\frac{QK^{T}}{\sqrt{d_{k}}})V$
We utilize the multi-head variation of attention where each head $H$ has the ability to generate its own attention weights:
(2) $\mathrm{MultiHead}(Q,K,V)=\mathrm{Concat}(\mathrm{h}_{1},\mathrm{h}_{2},\dots,\mathrm{h}_{n})$
The attention mechanism enables the model to adjust the weight given to different tokens in a sequence to optimally solve its objective. In LiDDA, we utilize a self-attention setup where $Q$, $K$, and $V$ are equvalently the marketing engagements performed by the member. We apply attention over the sequence $S$ and use the attention scores in the training classification task. The attention operation produces a representation of the sequence $E_{S}$ where $E_{S}\in\mathbb{R}^{N\times H\times{\text{dim}}}$. We aggregate the heads by averaging the scores across the head dimension to obtain $E^{\prime}_{S}\in\mathbb{R}^{N\times{\text{dim}}}$.
Different strategies exist for performing classification from attention scores. In BERT (Devlin et al., 2019), the authors created a special [CLS] token at the start and use only the learned representation of the first token. Other valid approaches include max-pooling or mean-pooling over the entire sequence. During inference, we compute the final attribution output as the summation of the attention weights across all heads and then normalize each row to 1.0 such that each row represents a path and each value in that row represents the $j$-th touchpoint in that sequence. We set masked values to 0 if a path is shorter than our maximum length.
In LiDDA, we keep all attention scores by flattening the attention output. The motivation behind this decision is to retain as much positional information as possible when computing the attention weights. Additionally, we limit our attention module to a single layer in order to reduce behavior where attention interpretability may be affected by abstract representations across timesteps as discussed in the next section.
2.4.1. Attention for Attribution
Prior works (li et al., 2018; Kumar et al., 2020) utilize attention-based crediting by applying a self-attention layer over an recurrent network. Our initial exploration of this approach found a heavy bias towards end-of-path touches. Using a bi-directional LSTM may help alleviate this effect, however we modified the approach to directly apply self attention over the input sequence. One issue with the LSTM encoder is that the hidden state contains information about previous timesteps, making attention interpretation less straight forward (Serrano and Smith, 2019). The topic of attention interpretability has been the subject of extensive research (Jain and Wallace, 2019; Wiegreffe and Pinter, 2019). The central concern is that the resulting weights may not faithfully align with the model final output. While our modification can still introduce interpretability issues in attention based models (Brunner et al., 2020), we observe within our domain that attention weights do impact model output. We share additional findings in Appendix B.2.
It is also possible to compute attribution based on the incremental propensity from the prior engagement (Shao and Li, 2011) or as an approximation of Shapely values (Yao et al., 2022). These methods are not necessarily superior. Counterfactual simulations can be computationally expensive and require accurately determining the counterfactual which is non-trivial. We did evaluate the attributions under the attention vs incremental propensity and found negligible differences (Table 1). We also observed that incremental propensity approach often produces early jumps causing front-loaded attribution as changes in propensity saturate with length.
Channel Incremental Attention
Channel A 14.4% 12.7%
Channel B 52.3% 53.4%
Channel C 29.5% 32.5%
Channel D 3.8% 1.5%
Table 1. Comparison of incremental propensity and attention weight from LiDDA model.
2.5. Imputation of Paid Media Touchpoints
LinkedIn’s outbound marketing channels are grouped into owned and paid channels. For owned channels, we can collect touchpoints from all actions at the user level, e.g., sends, opens, clicks, impressions, while for external channels, impression data at the user level are not available, e.g., due to privacy regulations like GDPR or CCPA (Appendix Figure 6).
Algorithm 1 Allocate Member Impressions
Inputs:
 External $\{I^{\rm ext}_{c,d},C^{\rm ext}_{c,d}\}$, clicks $J_{i,c,d}$
 Owned Impressions $\{I^{\rm own}_{\rm total},I^{\rm own}_{\rm clickers}\}$
 No-click paths $\mathcal{P}_{d}$, path‐weights $\{u_{k}\}$
Outputs: click‐imps $\{m_{j}\}$, path‐imps $\{I_{k}\}$
Step 1: Member‐only budget
$I_{d}=\sum_{c}I^{\rm ext}_{c,d}\;\frac{\sum_{i}|J_{i,c,d}|}{C^{\rm ext}_{c,d}}$
Step 2: Stochastic click allocation
$\alpha=\frac{I^{\rm own}_{\rm clickers}}{I^{\rm own}_{\rm total}},\quad r_{c,d}=\alpha\,\frac{I^{\rm ext}_{c,d}}{C^{\rm ext}_{c,d}}$
$m_{j}\;\sim\;\mathrm{Geom}\bigl{(}p=\tfrac{1}{r_{c(j),d}}\bigr{)},\quad L_{d}=I_{d}-\sum_{j}m_{j}$
Step 3: Stochastic leftover allocation
$p_{k}=\frac{u_{k}}{\sum_{\ell\in\mathcal{P}_{d}}u_{\ell}},\quad\lambda_{k}=L_{d}\,p_{k},\quad\tilde{I}_{k}\sim\mathrm{Poisson}(\lambda_{k})$
$I_{k}=\Bigl{\lfloor}\tilde{I}_{k}\,\frac{L_{d}}{\sum_{\ell}\tilde{I}_{\ell}}\Bigr{\rfloor}$
When a user clicks on an ad on an external platform they land on a LinkedIn microsite. We can use the URLs from microsites as a proxy for paid clicks at the user level. Meanwhile, we can also obtain daily aggregate campaign-level data, e.g, number of impressions or clicks for a campaign on a given day. Together with member-level clicks and impressions from owned channel, we can impute paid media impressions probabilistically.
Let $i$ index members, $d$ days, $c$ channels, $j\in J_{i,c,d}$ external‐click events, and $k\in\mathcal{P}_{d}$ no‐click paths (i.e., path without clicks from external channel). We perform the imputation in three steps (Algorithm 1): (i) Estimate member impression from external total impressions, yielding total ”budget” that needs allocation; (ii) Compute impression per clicks among the clickers $r_{c,d}$ and then draw impression from corresponding geometric distribution; (iii) Assign weights $u_{k}$ for no-click path $k$ based on campaign and location information, then draw Poisson counts of impressions for no-click path to allocate remaining impressions. Further information, including the step-by-step procedure, implementation notes, the equivalent convex‐optimization formulation, and its solution, is provided in Appendix A.1.
2.6. Sessionization
Another challenge in attribution modeling is that some channels generate far more touchpoints than others, creating a signal imbalance. For example, emails may be sent weekly, whereas in-feed ads can appear multiple times per day. Since our attribution paths have a fixed length, owned-channel touchpoints can dominate and crowd out paid (external) exposures. To address this, we apply two path-adjustment steps, sessionization and downsampling, to produce more balanced engagement sequences (Figure 2).
For sessionization, we group impressions from the same campaign on the same day into a single session‐level touchpoint for specific channels. For downsampling, we apply empirically derived, channel–action–specific drop rates to remove excess touchpoints. In both approaches, we recover the attribution credit to grouped/removed events according to the sampling probability. More details can be found in Appendix A.2
Figure 2. We group touchpoints from some owned channels, e.g., Linkedin Platform, within a session into a single representative touchpoint. The attribution given to this one touchpoint is distributed evenly among the grouped campaigns.
[…]
2.7. Output Calibration
MMM provides channel-level attribution and measures marketing incrementality in the context of macroeconomic factors but cannot trace individual member journeys. DDA captures those journeys at the interaction level but lacks certain signals (e.g. radio, events, TV, or privacy-restricted data). Two independent attribution systems may diverge, particularly if they utilize different interactions, and thus a unified approach is needed(Google, [n. d.]; Experimental, [n. d.]).
We introduce a training-time calibration that extract signals from our MMM model to our DDA model to form a unified macro level view for marketer, while still permitting fine-grained, path-level attributions. Let $a_{j}(\theta)$ be the DDA attribution weight for touchpoint $j$, $A^{\rm DDA}_{c}(\theta)=\sum_{j:\mathrm{chan}(j)=c}a_{j}(\theta)$ the total DDA weight for channel $c$, $A^{\rm MMM}_{c}$ the corresponding channel-level attribution from MMM, $\mathcal{L}_{\rm DDA}(\theta)$ the usual DDA training loss, and $\beta\geq 0$ a tuning parameter. We minimize the combined loss:
$\mathcal{L}_{\rm total}(\theta)=\mathcal{L}_{\rm DDA}(\theta)\;+\;\beta\sum_{c}\bigl{(}A^{\rm DDA}_{c}(\theta)-A^{\rm MMM}_{c}\bigr{)}^{2},$
and provide path level calibration in Algorithm 2.
This calibration “teaches” the DDA model, during training, to respect the macro-level attribution from MMM, eliminating the need for any post-hoc rescaling, while preserving detailed, campaign- and engagement-level attribution within each channel. Details on batch level calibration and path level calibration can be found in Appendix A.3
Algorithm 2 Train DDA Model with Path‐Level MMM Calibration
Inputs:
  • Paths dataset $\{(x_{i},y_{i},s_{i})\}_{i=1}^{N}$, where $x_{i}$ are features, $y_{i}$ labels, and $s_{i}$ the sequence of length $L_{i}$
  • MMM attribution $a_{c}^{\rm MMM}$, channel distribution $\pi_{c}^{\rm MMM}$
  • DDA network $g(\cdot;\theta)$ with parameters $\theta$, that output conversion probability $f$, and attention weight $a$.
  • Hyperparameters: calibration weight $\beta$, learning rate $\eta$, epochs $T$, batch size $B$
Output: Trained parameters $\theta$
for $\mathrm{epoch}=1$ to $T$ do
     Shuffle dataset
     for each batch $\mathcal{B}$ of size $B$ do $\triangleright$ Forward pass
         for each path $i\in\mathcal{B}$ do
              Compute DDA attributions $a_{i}(c;\theta)$ for $c=1,\dots,C$
              Compute path-level channel distribution $\pi^{(i)}_{c}$
              Compute path‐level IPW weights $w^{(i)}_{c}=\frac{\pi^{(i)}_{c}}{\pi_{c}^{\rm MMM}}$
              ”Recover” path-level MMM attribution as target
$\tilde{t}^{(i)}_{c}=w^{(i)}_{c}\,a_{c}^{\rm MMM},\quad t^{(i)}_{c}=\frac{\tilde{t}^{(i)}_{c}}{\sum_{d}\tilde{t}^{(i)}_{d}}$
         end for
         Compute base DDA loss
$\mathcal{L}_{\rm DDA}=\frac{1}{|\mathcal{B}|}\sum_{i\in\mathcal{B}}\ell\bigl{(}f_{i}(\cdot;\theta),y_{i}\bigr{)}$
         Compute calibration loss (MSE or KL):
$\mathcal{L}_{\rm calib}=\frac{1}{|\mathcal{B}|}\sum_{i\in\mathcal{B}}\sum_{c=1}^{C}\bigl{(}a_{i}(c;\theta)-t^{(i)}_{c}\bigr{)}^{2}$
         Total loss:
$\mathcal{L}=\mathcal{L}_{\rm DDA}\;+\;\beta\,\mathcal{L}_{\rm calib}$
$\triangleright$ Backward and update
         $\theta\leftarrow\theta-\eta\,\nabla_{\theta}\mathcal{L}$
     end for
end for
3. Experimentation Results
The goal of the attribution model is to proportionally quantify the impact of each marketing action towards a conversion. Since there is no ground-truth data on the contribution of each action towards a conversion, we perform the following offline and online validations to evaluate the attribution model.
3.1. Offline Validation
In offline validation, we seek ensure the model’s accuracy and stability as these are intuitively desirable properties of any attribution system. Because our offline process only has access to observational data, we primarily focus on typical ML performance as guardrails rather than actual attribution accuracy, which we will discuss in Section 3.2.
3.1.1. Prediction Stability
We first assess the model’s performance using binary classification metrics, aiming for high predictive accuracy. We hold 10% training data for validation and separate the rest into 10 subsets. We retrain an attribution model on each subset, evaluate it on the holdout dataset, and record its performance. The resulting prediction performances converge ($>0.97$ for ROC-AUC and PR-AUC), confirming that the attribution model has high prediction power and exhibits consistent results.
We also evaluate the stability of predicted conversion rate using the Anderson-Darling statistic (Stephens, 1974; Anderson and Darling, 1952). The result shows that the predicted conversion rate are consistent over time. Further information is provided in Appendix B.1.
3.1.2. Attribution Stability
We perform a similar check on our model’s stability in producing attribution results. For each action, we bootstrap sample that action’s weights with replacement for 100 times, calculate the mean of the samples, and analyze the distribution of the 100 means to evaluate the stability of the attribution weights (Figure 3). We observe that the mean weights of each action converge. While this approach does not have an intrinsic absolute threshold, we use these values as guardrails to monitor for potential issues.
Figure 3. Distribution of bias in the bootstrap means of the action weights, calculated as the difference between each bootstrap mean and the actual sample mean. The bias distributions center around 0.
[…]
3.1.3. Ablation on Model Input
We conducted an ablation study to quantify the impact of each model component by training and evaluating DDA model with particular features removed, including member/company embeddings, campaign embeddings, date embeddings, or all embeddings (sequence only). We compared ROC-AUC against the baseline and measured how the channel-level attribution mix deviated (mean squared deviation on percentage) from the full model. Removing campaign or date features caused small AUC drops (–0.1% and –0.3%), while omitting entity embeddings and using sequence alone reduced AUC more substantially (–1.7% and –5.1%). In terms of attribution stability, date removal produced the largest shift (MSD=49.3), with campaign and entity removals and the sequence-only variant showing smaller but non-negligible deviations (MSD=8.5, 14.4, and 11.8, respectively). Details are in Appendix B.3.
3.2. Online Validation
Offline validation does not give us insight into whether the attention weight based attribution are close to actual member behavior. Thus, we run experiments for online validations to check how close our model’s attributions align with real-world marketing efforts. We can either run member level test for for owned channels or geo-level experiments for external paid channels. We conduct incremental lift tests for email channel. In particular, we randomly split members into two groups. We select certain campaigns ($\mathcal{T}^{h}$) to hold out email sending during the experiment in the control group and conduct business as usual in the treatment group. It is important to note that email sending was sparse across both members and time, limited to approximately one email per week to a small set of selected members within the connection graph. Additionally, AA tests were conducted as a basic validation of the experimental setup, and no statistically significant differences in conversion rates were found between groups.
To validate member-level experiment, we simulate a holdout experiment within the treatment group by creating counterfactual conversion paths, where the holdout campaign touchpoints ($t\in\mathcal{T}^{h}$) are omitted. This approach helps estimate what the conversion path would look like had the member been in the control group. Define two conversion probabilities: $P(Y=1|Z=1,S=S_{1})$ with original path $S_{1}$, and $P_{1}(Y=1|Z=1,S=S_{1}-\mathcal{T}^{h})$ with counterfactual path $S_{1}-\mathcal{T}^{h}$, where $Z=1$ index the treatment group. The corresponding attribution of touchpoints $t\in\mathcal{T}^{h}$ can be defined as:
(3) $a({\mathcal{T}^{h}})=\frac{P_{1}(Y=1|S=S_{1})-P_{1}(Y=1|S=S_{1}-\mathcal{T}^{h})}{P_{1}(Y=1|S=S_{1})}$
where, $P_{1}(\cdot)=P(\cdot|Z=1)$. We compare the distribution of DDA attribution against the measured attribution from experiment.
A similar formulation can be constructed using $P_{0}(\cdot)$. We didn’t go with this route for validation, as the simulation of counterfactual paths ($S_{0}+\mathcal{T}^{h}$) by adding touches requires more assumptions such as sequencing and campaigns which may introduce bias and noise.
3.2.1. Inverse Propensity Weighting
From the experimentation data, we can calculate $\hat{P}_{1}(Y=1|S=S_{1})$ and $\hat{P}_{0}(Y=1|S=S_{1}-\mathcal{T}^{h})$ easily, by simply counting percentage of conversion in the treatment and control group. In order to obtain estimate of $P_{1}(Y=1|S=S_{1}-\mathcal{T}^{h})$, we use inverse propensity weighting (IPW)(Hernán and Robins, 2006). Denote $e_{i}$ as the propensity score for member $i$ and $Z_{i}$ as the indicator for member $i$ receiving campaigns as in the treatment group (=1) and for being in the control group (=0). Denote $E_{M_{i}}$ as the member embeddings for member $i$ that is computed by LinkedIn.
Step 1: Estimate each member’s treatment propensity, and corresponding IPW
(4) $e_{i}=P(Z_{i}=1|E_{M_{i}}),\qquad w_{i}=\begin{cases}1,&\text{if }Z_{i}=1\\ \frac{e_{i}}{1-e_{i}},&\text{if }Z_{i}=0\end{cases}$
Step 2: Compute the attribution
(5) $\hat{a}^{\exp}(\mathcal{T}^{h})=\frac{\hat{P}_{1}(Y=1|S=S_{1})-\hat{P}_{1}(Y=1|S=S_{1}-\mathcal{T}^{h})}{\hat{P}_{1}(Y=1|S=S_{1})}$
by plugging in $\hat{P}_{1}(Y=1|S=S_{1})=(\sum_{i:Z_{i}=1}Y_{i})/(\sum_{i:Z_{i}=1}1)$, and
(6) $\hat{P}_{1}(Y=1|S=S_{1}-\mathcal{T}^{h})=\sum_{i:Z_{i}=0}\frac{w_{i}Y_{i}}{\sum_{i:Z_{i}=0}w_{i}}.$
For validation, we then compute the attention based attribution, i.e.,
(7) $\hat{a}^{DDA}(\mathcal{T}^{h})=\frac{\sum_{i:Z_{i}=1}\sum_{t\in\mathcal{T}^{h}}a_{i}^{DDA}(t)}{\sum_{i:Z_{i}=1}1}$
where $a_{i}^{DDA}(t)$ is attention based attribution for member $i$ for touchpoint $t$. Further details, including validation of IPW and two additional attribution measures are available in Appendix B.4.
3.2.2. Experiment Results
We chose three incrementality tests in the email channel to validate the attribution model and the primary metric is attribution defined in eq. [B.4]. Our goal is to evaluate the effectiveness of the acquisition email series. The workflow is illustrated in Figure 4 and the comparisons are listed in Table 2.
Figure 4. The online validation refers to the comparison between measured lift from experiment data and the simulated lift output by LiDDA and the aforementioned calibration
[…]
Test index $\hat{a}^{\exp}$ $\hat{a}^{DDA}$ $\hat{d}^{DDA}$
1 9.68% (1.47%, 20.02%) 7.65% (4.95%, 10.07%) -2.23% (-5.85%, 0.06%)
2 14.05% (3.35%, 24.75%) 14.96% (11.78%, 18.24%) 0.95% (-0.89%, 2.25%)
3 12.54% (2.10%, 27.69%) 15.21% (11.65%, 19.34%) 2.41% (-0.38%, 5.05%)
Table 2. Experimental validation for 3 incrementality tests. Each cell contains the point estimate and its corresponding 95% confidence intervals.
We compare DDA generated attribution $\hat{a}^{DDA}$ with ”ground truth” $\hat{a}^{\exp}$ and compute their difference, i.e. $\hat{d}^{DDA}=\hat{a}^{DDA}-\hat{a}^{\exp}$.
Confidence intervals of $\hat{a}^{\exp}$, $\hat{a}^{DDA}$, and $\hat{d}^{DDA}$ were derived using bootstrap sampling. We notice that the gaps between $\hat{a}^{\exp}$ and $\hat{a}^{DDA}$ are not statistically significant. We concluded that our DDA attention based attribution align with the experimental result.
4. Application to LinkedIn Ads Measurement
The LiDDA model architecture was first applied to LinkedIn’s GTM marketing. We have since integrated a variant of LiDDA into LinkedIn’s Ad Platform to provide advertisers with comprehensive funnel insights through LinkedIn Ads. Previously, measurement of ad impact on the Ads platform heavily relied on last-touch attribution, which often undervalues upper-funnel ads, potentially resulting in suboptimal budget allocation. In contrast, data-driven attribution distributes credit across multiple ad interactions throughout the buyer’s journey, accounting for the impact of upper-funnel advertising efforts on business outcomes.
When applying LiDDA to the LinkedIn Ads use case, we also supply information about the advertiser, via a company embedding for the advertiser. This is concatenated with embeddings representing the member and their company. Similar to our GTM marketing model, we also include a touchpoint sequence representation (encoding the sequence of touch types, e.g., ”Onsite lead generation sponsored status update video ad impression”), touchpoint embeddings derived from campaign text, and positional encodings (representing the time between each touchpoint and the conversion). While GTM marketing has access to touchpoints across multiple channels, the Ads platform currently only has access to LinkedIn touchpoints. When applying the model to lead conversions obtained through LinkedIn’s Lead Generation form, on average the attribution weights show a time decay effect: touches closer to the conversion tend to get more credit (Figure 5). Meanwhile, We also observe the weight assigned to a specific touchpoint can vary, influenced by factors like the objective type, ad format, interaction type (impression, click), ad creative, days before conversion. We surface these insights at both path‐level detail and aggregated campaign/creative reports (Details in Appendix B.5).
Figure 5. Attribution weights for LinkedIn lead conversions show a time decay effect in aggregate; ad touches closer to the conversion tend to get more credit.
5. Conclusion
In this paper, we introduce LiDDA, a unified transformer-based data-driven attribution approach we developed and deployed at LinkedIn. We demonstrated the innovations to DDA methodology that combines an attention mechanism, positional embeddings, touchpoint imputation, and MMM calibration. We also shared learnings from implementation and deployment that are broadly applicable to the Marketing and Ad tech fields.
Appendix A Additional Details on Methods
A.1. Three‐Step Member Impression Allocation
Let $d$ index days, $c$ index paid channels, $j\in J_{i,c,d}$ index click‐events for member $i$, and $k\in\mathcal{P}_{d}$ index no‐click paths on day $d$. We observe external aggregates $(I^{\rm ext}_{c,d},\,C^{\rm ext}_{c,d})$, owned‐channel totals $(I^{\rm own}_{\rm total},C^{\rm own}_{\rm total},I^{\rm own}_{\rm clickers},C^{\rm own}_{\rm clickers})$, and a nonnegative weight $u_{k}$ for each no‐click path.
A.1.1. Member‐Only Impression Budget
Compute the number of impressions on channel $c$ and day $d$ attributable to known members by scaling the external impressions by the fraction of clicks from members:
$C^{\rm mem}_{c,d}=\sum_{i}\lvert J_{i,c,d}\rvert,\qquad I^{\rm mem}_{c,d}=I^{\rm ext}_{c,d}\,\frac{C^{\rm mem}_{c,d}}{C^{\rm ext}_{c,d}}.$
Summing over channels gives the total daily member‐impression budget:
$I_{d}\;=\;\sum_{c}I^{\rm mem}_{c,d}.$
A.1.2. Impute and Pre‐Allocate to Clicks
First, derive a down‐weight factor $\alpha<1$ from owned‐channel behavior:
$R^{\rm own}_{\rm all}=\frac{I^{\rm own}_{\rm total}}{C^{\rm own}_{\rm total}},\quad R^{\rm own}_{\rm clickers}=\frac{I^{\rm own}_{\rm clickers}}{C^{\rm own}_{\rm clickers}},\quad\alpha=\frac{R^{\rm own}_{\rm clickers}}{R^{\rm own}_{\rm all}}.$
Apply $\alpha$ to the raw external impressions‐per‐click to obtain the clickers’ expected impressions per click on $(c,d)$:
$R^{\rm ext}_{\rm all}=\frac{I^{\rm ext}_{c,d}}{C^{\rm ext}_{c,d}},\qquad r_{c,d}=\alpha\,R^{\rm ext}_{\rm all}\;\approx\;\mathbb{E}\bigl{[}\#\mathrm{imps}\mid\mathrm{click},c,d\bigr{]}.$
Then allocate to each click $j\in J_{i,c,d}$ at least one impression by either
[…]
and compute the pre‐allocated sum and leftover:
$I^{\rm pre}_{d}=\sum_{j}m_{j},\qquad L_{d}=I_{d}-I^{\rm pre}_{d}.$
A.1.3. Probabilistic Allocation of Leftover
Assign the remaining $L_{d}$ impressions across no‐click paths $k\in\mathcal{P}_{d}$ proportionally to weights $u_{k}$. Define
$p_{k}=\frac{u_{k}}{\sum_{\ell\in\mathcal{P}_{d}}u_{\ell}},\qquad\lambda_{k}=L_{d}\,p_{k}.$
Draw $\tilde{I}_{k}\sim\mathrm{Poisson}(\lambda_{k})$ independently, then rescale to exactly exhaust $L_{d}$:
$I_{k}=\Bigl{\lfloor}\tilde{I}_{k}\times\frac{L_{d}}{\sum_{\ell}\tilde{I}_{\ell}}\Bigr{\rfloor},\qquad\sum_{k}I_{k}=L_{d}.$
A.1.4. Result
Each click‐event $j$ receives $m_{j}\geq 1$ (on average $r_{c,d}$), each no‐click path $k$ receives $I_{k}\geq 0$, and
$\sum_{j}m_{j}\;+\;\sum_{k}I_{k}\;=\;I_{d},$
so the member‐impression budget is exactly allocated.
A.1.5. Remarks
We note that:
  • Both Geometric distribution and Poisson distribution can be appoximated by Normal distribution if needed;
  • $u_{k}\propto\mathbb{E}\bigl{[}\#\mathrm{imps}\mid\mathrm{noClick},k\bigr{]}$ can be assigned by contextual information such as corresponding campaign audience and location;
  • $r_{c,d}$ can be estimated with more complex methods, e.g., $r_{i,c,d}=f(x_{i},x_{c},x_{d})$ via a NN that leverage both member level data from owned channel and aggregate data from external channels.
A.1.6. Joint Convex‐Optimization Formulation
Instead of treating click allocation (Step 2) and leftover allocation (Step 3) separately, we can solve them in one convex‐optimization step.
Introduce: (i) Variables $x_{j}$ for the impressions assigned to click‐event $j\in J_{*,*,d}$; (ii) Variables $y_{k}$ for the impressions assigned to no‐click path $k\in\mathcal{P}_{d}$; (iii) Total budget $I_{d}$ from Step 1; (iv) Positive weights
[…]
reflecting the expected imps‐per‐click and the no‐click path priorities.
We then solve the following optimization:
$\underset{\{x_{j}\},\{y_{k}\}}{\text{maximize}}$ $\sum_{j\in J_{*,*,d}}w_{j}\,\log(x_{j})\;+\;\sum_{k\in\mathcal{P}_{d}}u_{k}\,\log(y_{k})$
subject to $\sum_{j\in J_{*,*,d}}x_{j}\;+\;\sum_{k\in\mathcal{P}_{d}}y_{k}\;=\;I_{d},$
[…]
KKT‐based solution. The Karush–Kuhn–Tucker conditions imply that at optimum there is a multiplier $\lambda>0$ such that
$x_{j}=\max\!\Bigl{(}1,\;\frac{w_{j}}{\lambda}\Bigr{)},\qquad y_{k}=\frac{u_{k}}{\lambda},$
and $\lambda$ must satisfy:
(8) $\sum_{j}\max\!\Bigl{(}1,\;\frac{w_{j}}{\lambda}\Bigr{)}+\sum_{k}\frac{u_{k}}{\lambda}=I_{d}.$
Eq. 8 can be solved by Algorithm 3, with complexity $O(J\log(J))$.
To summarize, the solution:
1. Honors the total budget: $\sum_{j}x_{j}+\sum_{k}y_{k}=I_{d}$.
2. Guarantees each click‐bucket $x_{j}\geq 1$ whenever $w_{j}>0$.
3. Allocates impressions in proportion to the expected click demand $w_{j}$ and the no‐click weights $u_{k}$.
By choosing $w_{j}=r_{c,d}$ and $u_{k}$ as in Steps 2–3, we recover a deterministic analogue of the previous stochastic Geometric+Poisson draws.
Algorithm 3 Compute Optimal $\lambda$ and Allocations
Input: Click‐weights $\{w_{j}\}_{j=1}^{J}$, path‐weights $\{u_{k}\}_{k=1}^{K}$, total budget $I_{d}$
Output: Impressions $\{x_{j}\},\{y_{k}\}$
Sort $w_{(1)}\geq w_{(2)}\geq\dots\geq w_{(J)}$
Compute $S_{u}\leftarrow\sum_{k=1}^{K}u_{k}$
$S_{w}\leftarrow\sum_{j=1}^{J}w_{(j)}$
for $m=0$ to $J$ do
     $S_{w}^{(m)}\leftarrow S_{w}-\sum_{j=1}^{m}w_{(j)}$
     $\lambda_{m}\leftarrow\frac{S_{w}^{(m)}+S_{u}}{\,I_{d}-m\,}$
end for
Find the unique $m^{*}$ such that
$w_{(m^{*})}\;\leq\;\lambda_{m^{*}}\quad\text{and}\quad w_{(m^{*}+1)}\;>\;\lambda_{m^{*}},$
with conventions $w_{(0)}=\infty$, $w_{(J+1)}=0$.
$\lambda\leftarrow\lambda_{m^{*}}$
for each click $j=1,\dots,J$ do
     $x_{j}\leftarrow\max\bigl{(}1,\;w_{j}/\lambda\bigr{)}$
end for
for each path $k=1,\dots,K$ do
     $y_{k}\leftarrow u_{k}/\lambda$
end for
Figure 6. The top row shows an original path with all impressions and clicks a member has interacted with prior to making a conversion. We cannot collect any paid media impressions due to privacy regulations like GDPR and CCPA, and we use a fixed length for paths to use in modeling. The middle row shows the same path after these interactions are dropped. We impute probabilistic impressions into the paths and group impressions from owned channels, LinkedIn Platform as shown in the bottom row for modeling.
[…]
Figure 7. Probabilistic touchpoints are imputed in paths in two steps. First, we add impressions to existing proxy-click events based on impression to click ratio of each channel. Second, we distribute the remaining impressions across members online each day.
[…]
A.2. Path Pre-processing
A.2.1. Sessionization
We group impressions from certain channels in a path as a single touchpoint within a session if they are from the same campaign on the same day, as detailed in Algorithm 4. This helps reduce noise, particularly from on-platform channels where a member may observe the same campaign repeatedly due to scrolling behavior. With this approach, we model a single event however we retain the downsampling links for post-model processing to reallocate credit equally across the grouped events.
A.2.2. Downsampling
We further address signal bias by adjusting the retention rate of specific channel, action combinations according to tunable downsampling rates. We obtain these rates using an independent analysis to estimate the comparable impact of each channel using signals such as spend and outcomes. With these ratios, we randomly ”drop” some touchpoints to yield the expected ratio as described in Algorithm 5. We observe that this pre-processing step yields more reliable attribution results and path compositions better aligned with aggregate data.
Our attribution system is complex and heterogeneous, as a result, certain heuristic is necessary. Meanwhile, we perform comprehensive evaluation on these heuristics. For example, to downsample owned channel activities, we estimate a scaling factor to bring the number of actions, clicks/impressions, to the same order of magnitude of actions from other channels. And we evaluate the impact through offline and online experiment.
Algorithm 4 Sessionization methodology
Member $path$ with marketing $touchpoint$ metadata
for $path$ do
     Group all impressions by $campaign$, $channel$, and $date$
     for $group$ do
         Select most recent $touchpoint$ by timestamp
         Record array of remaining $touchpoint$
         Store array in $touchpoint$ metadata
     end for
     Output truncated $path$ with linkage metadata
end for
Algorithm 5 Downsampling methodology
Member $path$ with marketing $touchpoint$ metadata
Sampling $ratio$ per $action$ and $channel$ pair
for $path$ do
     Group all touchpoints by $action$ and $channel$
     for $group$ do
         Sort $touchpoints$ by timestamp
         Compute number of $touchpoints$ $N$ based on $ratio$
         Retain $N$ latest $touchpoints$
     end for
     Output truncated $path$
end for
A.3. Training Time Calibration
A.3.1. Batch‐Level Calibration
Let a training batch consist of $N$ paths. Denote by
$a^{\rm DDA}_{i}(c;\theta)=\sum_{t\in c}a_{i}^{\text{DDA}}(t;\theta),\quad(i=1,\dots,N,\;c=1,\dots,C)$
the DDA model’s attribution share for channel $c$ on path $i$. Define the batch‐aggregated channel totals
$A^{\rm DDA}_{c}(\theta)=\frac{1}{N}\sum_{i=1}^{N}a^{\rm DDA}_{i}(c;\theta),\qquad A^{\rm MMM}_{c}=a^{\rm MMM}_{c}$
where $a^{\rm MMM}_{c}$ is the global MMM‐derived share for channel $c$. We then add one of the following batch‐level penalties:
Mean‐Squared Error (MSE) Penalty
$\mathcal{L}_{\rm calib}^{\rm batch}=\sum_{c=1}^{C}\Bigl{(}A^{\rm DDA}_{c}(\theta)-A^{\rm MMM}_{c}\Bigr{)}^{2};$
Kullback–Leibler (KL) Penalty
$\mathcal{L}_{\rm calib}^{\rm batch}=\sum_{c=1}^{C}A^{\rm MMM}_{c}\,\ln\!\frac{A^{\rm MMM}_{c}}{A^{\rm DDA}_{c}(\theta)}.$
These terms are computed once per batch and combined with the base DDA loss:
$\mathcal{L}_{\rm total}=\mathcal{L}_{\rm DDA}(\theta)\;+\;\beta\,\mathcal{L}_{\rm calib}^{\rm batch}.$
Backpropagation propagates gradients from the channel‐sum residuals into each $a^{\rm DDA}_{i}(c;\theta)$.
A.3.2. Path‐Level Calibration
For each path $i$, let $\pi^{(i)}_{c}$ be the empirical fraction of touches in channel $c$:
$\pi^{(i)}_{c}=\frac{\#\{\text{touchpoints of }c\text{ in path }i\}}{\text{length}(i)}.$
Let $\pi^{\rm MMM}_{c}$ be the probability of any touchpoint’s channel being $c$ (i.e., estimate from all paths), and $a_{c}^{MMM}$ be the ”attribution” of $c$ from MMM. Define a per‐path, per‐channel weight
$w^{(i)}_{c}=\frac{\pi^{(i)}_{c}}{\pi^{\rm MMM}_{c}}.$
Form the unnormalized target
$\tilde{t}^{(i)}_{c}=w^{(i)}_{c}\;a^{\rm MMM}_{c},$
and then normalize to ”recover” a valid path-level MMM attribution
$t^{(i)}_{c}=\frac{\tilde{t}^{(i)}_{c}}{\sum_{d=1}^{C}\tilde{t}^{(i)}_{d}},\quad\sum_{c=1}^{C}t^{(i)}_{c}=1.$
We then penalize the divergence between DDA’s prediction and this target for each path:
Per‐Path MSE
$\mathcal{L}_{\rm calib}^{\rm path}=\sum_{c=1}^{C}\Bigl{(}a^{\rm DDA}_{i}(c;\theta)-t^{(i)}_{c}\Bigr{)}^{2}.$
Per‐Path KL
$\mathcal{L}_{\rm calib}^{\rm path}=\sum_{c=1}^{C}t^{(i)}_{c}\,\ln\!\frac{t^{(i)}_{c}}{a^{\rm DDA}_{i}(c;\theta)}.$
As before, the total loss is
$\mathcal{L}_{\rm total}=\mathcal{L}_{\rm DDA}(\theta)\;+\;\beta\,\mathcal{L}_{\rm calib}^{\rm path},$
and gradients flow from each path‐level residual into the DDA model parameters $\theta$.
Appendix B Additional Offline Analysis
B.1. Model Stability
We provide visualization of model stability on prediction and attribution in Figure 8, Figure 9 and Figure 3.
Figure 8. The ROC-AUC and PR-AUC of the attributions models retrained on 10 different subsets of training data and evaluated on the same holdout dataset. The models’ performances converge.
Figure 9. Anderson-Darling statistic comparing the distribution of conversion propensities each day to the distribution from the day a week ago. Days where the propensity distributions are significantly different are marked in red.
[…]
B.2. Attention Interpretability
When using attention weights in attribution, we would expect that the computed weights quantify the contribution a touchpoint has towards the prediction of a conversion outcome. Therefore, if we were to modify the weight distribution we should observe a change in the predicted outcome. In this test, we train two models on the same dataset. The baseline model is trained using the standard attention setup described in this paper. The modified model randomly permutes the attention weights at a path-level during inference forward passes.
We train both models for 10 epochs each using identical batch sizes and hyperparameters. After training, we perform an inference pass on a shared holdout dataset and measure ROC-AUC. For the baseline model, we perform a single inference pass. For the modified model, we perform 10 inference passes to capture the variance introduced by the weight permutation process. We repeat this entire process 30 times for each model.
In Figure 10, we observe that the baseline model generally yields higher performance with a median ROC-AUC of 0.956 while the modified model has a median ROC-AUC value of 0.928 (-2.93%). While this is not a substantial change, we observe a larger negative impact on the consistency of results. The AUCs of both models are noted to be high. There are two explainations for this. One reason is due to resampling of the data creating an artificially high rate of conversions. The model is measured on uncalibrated results. Secondly, we note that in our domain, converting paths have a notably different composition which facilities an easier classification task.
Figure 10. The plot compares the ROC-AUC against a common holdout for both the baseline and modified attention models in our study. The baseline model consists of 30 datapoints with 1 validation step per trial. The modified model consists of 300 datapoints with 10 validation steps per trial.
In Figure 11, we plot the results of the 30 modified trials sorted in median ascending order. 23 of the 30 trials ranked below the 20th percentile compared to the baseline model. We conclude from this test that the introduction of weight noise has a material impact on model performance and therefore provides some explanability power.
Figure 11. This plot shows the per-trial results for the modified model. The trials are sorted in ascending order according to their median ROC-AUC value. We observe that 23 of the 30 models fall below the bottom 20th percentile of the baseline trials.
B.3. Ablation
B.3.1. Methodology
We performed an ablation study to estimate the contribution of different components of our model introduced in this paper. We generated fixed training, validation, and inference datasets to use across each simulation where a new model was trained and inferenced with the indicated component removed. In each simulation, the removal of the component was performed by modifying the code to omit the component from the graph. All hyperparameters and were held constant, and each model was trained for 5 epochs.
After training, we evaluated the ROC-AUC metrics against the validation dataset which contains a resampled population ratio of non-converting to converting of 10:1. We also evaluated the attribution weight distribution across channels for each of the models.
We tested the following components of the model:
  • Baseline: The unmodified model as described in the paper
  • No-Entity: Member and company embedding inputs removed
  • No-Campaign: Touchpoint campaign embedding inputs removed
  • No-Date: Touchpoint date embedding inputs removed
  • Seq-Only: All embeddings removed leaving only the sequence and positional encoding
B.3.2. Classification Results
We first evaluated the ablation effects on the final classification power of the model by comparing the predicted model score to . We note high nominal AUC values as the simulations were performed on a resampled dataset. We observe that there is a moderate performance loss for the touchpoint level date and campaign features. Removal of entity embeddings had a larger effect, and removal of all components had the most negative effect. We also note that utilizing the sequence only allows for substantial predictive power given that converting members in our domain behave differently than non-converting members.
Version ROC-AUC % Change
1. Baseline 0.990
2. No-Campaign 0.989 -0.1%
3. No-Date 0.984 -0.3%
4. No-Entity 0.973 -1.7%
5. Seq-Only 0.939 -5.1%
Table 3. ROC-AUC performance measured on a constant validation dataset after 5 epochs of training. Percentage change vs. the baseline model is shown.
B.3.3. Weight Results
We next evaluated the impact of the component removal on the measured attribution credit across a grouping of our four marketing channel families. We compute the the mean squared deviation in total attribution from the baseline model across the four channels as a measure of volatility in the ablation. We observe that all ablations result in some changes to attribution however removing the dates alone causes the most significant shifts.
Version Mean Squared Deviation
1. Baseline
2. No-Campaign 8.5
3. No-Date 49.3
4. No-Entity 14.4
5. Seq-Only 11.8
Table 4. The percent attribution allocated across four marketing channels is computed using the indicated LiDDA model version. The absolute deviation for each channel vs. the baseline model is computed. The reported number is the mean squared deviation for that model, across all channels, compared with the original basline mix.
B.4. Recovering Distribution
The inverse propensity score weight can help balance the converted members in the treatment group and members in control group for fairer comparison between model predicted conversions and observed conversions from online experiments. In Figure  12, we create weight-scaled pseudo-populations through $w_{i}$ for those who converted in the treatment group and for those in the control group. We observe that IPW functions as expected to adjust the population from converted member in treatment group and control group to a closer match.
Figure 12. The top chart shows a mismatch between conversion propensity score distribution for converted members in the treatment group vs. control members. The bottom chart shows that conversion propensity score distributions are better matched in the pseudo-populations where we replicate the members through $w_{i}$ to balance the two populations.
[…]
Without IPW, we can calculate a raw estimate of $a^{\text{raw}}({\mathcal{T}^{h}})$, which can be biased.
$a^{\text{raw}}({\mathcal{T}^{h}})=\frac{\hat{P}_{1}(Y=1|S=S_{1})-\hat{P}_{0}(Y=1|S=S_{1}-\mathcal{T}^{h})}{\hat{P}_{1}(Y=1|S=S_{1})}.$
Test index $a^{\text{raw}}$ $\hat{a}^{DDA}$ $\tilde{a}^{DDA}$
1 8.72% (0.01%, 16.11%) 7.65% (4.95%, 10.07%) 12.54% (10.52%, 14.73%)
2 9.73% (1.71%, 17.75%) 14.96% (11.78%, 18.24%) 19.61% (16.81%, 23.44%)
3 16.65% (3.93%, 29.37%) 15.21% (11.65%, 19.34%) 9.54% (7.53%, 11.75%)
Table 5. Additional comparison for 3 incrementality tests. Each cell contains estimated attribution and its corresponding 95% confidence intervals.
We also calculate a plug-in estimate $\tilde{a}^{DDA}$ of the attribution with DDA model, similar as eq. [5]. Let $\pi(S)$ be the predicted conversion for touchpoint sequence $S$ from DDA model, we can compute $\tilde{P}_{1}(Y=1|S=S_{1})=(\sum_{i:Z_{i}=1}\pi(S_{1,i}))/(\sum_{i:Z_{i}=1}1)$, and $\tilde{P}_{1}(Y=1|S=S_{1}-\mathcal{T}^{h})=(\sum_{i:Z_{i}=1}\pi(S_{1,i}-\mathcal{T}^{h}))/(\sum_{i:Z_{i}=1}1)$. We compared these two estimate with ”ground truth” $\hat{a}^{\exp}$ in Table 5.
B.5. Application to Ad Platform
Touchpoint-level attribution from LiDDA provides very granular insights, where we see varying credit assignment to earlier touchpoints. In the provided examples in Figure 13, touches closer to the conversion tend to have higher weights; however, the weight assigned to a specific touchpoint can vary, influenced by factors like the objective type, ad format, interaction type (impression, click), ad creative, days before conversion, etc. Using these types of signals, the model can learn to allocate credit across all touchpoints in the converting paths. This data can then be surfaced to advertisers on LinkedIn’s Ad platform at different levels of aggregation, in a variety of ways, including in the Campaign Manager Reporting table (i.e., showing the number of conversions attributed to each campaign, creative, etc.), and in a plot in Measurement Insights (showing a time series of conversions attributed to different categories of campaigns).
Figure 13. The attribution weights extracted from the model can provide granular insights. Here we show example attribution weights for two LinkedIn lead conversions, as a function of touches until conversion (x-axis) and touch type (hue). While in aggregate we see a time-decay effect (i.e., touches closer to the conversion tend to get more credit), the credit allocated to each touch in a given path can vary. For example, in the bottom path, the initial blue touches receive relatively more credit than the next few touches, and the red ”click” three touches before the conversion receives the highest credit.